Tina earned $1.38 per day for two days as interest on her $1,000.00 investment. Select all the expressions you could use to find the percent interest she earned in total.
A) 2.76•1000/100
B) 2(1.38)•1000
C) 1.38/1000
D) 2(1.38)/1000
E) 2.76/1000

Answer by Theo(13342)

she earned two * 1.38 interest for the 2 days.
that would be equal to 2.76.
the percent interest rate on that would be 2.76 / 1000 * 100
the interest rate is 2.76 / 1000 which is equal to .00276.
that is multiplied by 100 to to get .276%.
selection A gets you 2.76 * 1000 / 100 = = 276.
selection B gets you 2 * 1.38 * 1000 = 2.76 * 1000 = 276000.
selection C gets you 1.38 / 1000 = .00138
selection D gets you 2 * 1.38 / 1000 = 2.76 / 1000 = .00276.
selection E gets you 2.76 / 1000 = .00276.
if i understood the problem correctly, none of those answer is correct.
the answer should be 2 * 1.38 / 1000 = interest rate of .00276.
that need to be multiplied by 100 to get interest rate = .276%.
selection D and E will give you the correct interest rate, but they will not give you the correct interest rate percent, unless you multiply them by 100.
if selection A was 2.76 / 1000 * 100, or 2.76 * 100 / 1000, then it would be correct, but it's not.